OpenStudy (anonymous):

is x+0=0 always true?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (debbieg):

try a few values for x. see what shakes out. :)

OpenStudy (skullpatrol):

I think you mean x + 0 = x

OpenStudy (anonymous):

no thats what it says

OpenStudy (debbieg):

That would make a big difference, huh? ;) lol

OpenStudy (debbieg):

Ok, then again - try a few values for x. Plug them in. Is it true?

OpenStudy (debbieg):

You can't PROVE that it's ALWAYS true by plugging in values.... but you MIGHT be able to DISPROVE something by plugging in values.

OpenStudy (anonymous):

thats not what theyre asking the options are always trus, sometimes true, or never true

OpenStudy (debbieg):

I get what the question is. I'm trying to help you figure it out.

OpenStudy (anonymous):

thanks

OpenStudy (debbieg):

For example, if you can think of an x for which its true, and one for which its false, then you have your answer.

OpenStudy (anonymous):

okay so ive been taught if you subtract the zero from both sides than x equal 1

OpenStudy (debbieg):

No, I'm certain you haven't been taught that. And there is no need to "subtract the 0" from both sides, it's 0. subtracting it doesn't accomplish anything. Try some values for x, and see if the statement is true or false. Maybe =1, x=2 and x=0.

OpenStudy (anonymous):

the question states " is x+0=0 always, sometimes, or never true?" there is not much you can do with that.

OpenStudy (anonymous):

i dont understand what you are trying to tell me to do.

OpenStudy (phi):

is 1+0 = 0 true ?

OpenStudy (anonymous):

omg thank you soo much. God bless you :*

OpenStudy (phi):

what about is 0+0= 0 true ?

OpenStudy (anonymous):

so its sometimes true ?

OpenStudy (phi):

if x is 1 is x+0=0 true if x is 0 is x +0 =0 true ?

OpenStudy (phi):

sometimes sounds correct. there are some x's that don't work (most of them!) but there is 1 x (when x is 0) that does make it true.

OpenStudy (anonymous):

oh thank you so much again

OpenStudy (debbieg):

Sorry I disappeared @angel1234567, I had to get dinner on the table. @phi went exactly where I was trying to go. When it says "always, sometimes, or never true" what it really means is, "is this statement true for EVERY value of x, for SOME but not all values of x, or for NO values of x?" Remember, x is just a "placeholder" for a number. It isn't asking you to "solve" the equation - although you CAN solve it, simply by simplifying the "x + 0" expression on the left to "x": then you have x=0, which IS a solution, so it is at least ONE value for x that makes it a true statement. But as you can quickly see, by plugging in other values for x, like x=1 or x=12 or x=6000, it ISN'T ALWAYS true. So if it's true for at least one value for x, but not for all, then it is "sometimes true". :)